# Minimizing Human Intervention in Online Classification

### v1

@@ E.2 Asymptotic Lower bound on the Regret of CHC
Now for any $0<a\leq 1$, $\lambda(\{y\in[0,1]^{d},\prod y_{i}\leq a\})=a\sum_{k=0}^{d-1}\log(1/a)^{k}/k!$ (see Bárány and Larman, (1988), p.288). Consequently, $w(t)\leq 2t\sum_{k=0}^{d-1}\log(1/(2t))^{k}/k!$ whenever $0<t\leq 1/2$. Applying this to $t=1/n$ yields the result. ∎
When $\nu$ is uniform on some arbitrary polytope $P$, explicitly bounding the volume of the wet part of $P$ becomes significantly more difficult. Schütt, (1991) proved in their Lemma 1.8 that for some $t_{d}$ and $c_{d}$ that only depends on $d$, if $0<t<t_{d}$,
@@
A natural idea to generalize Theorem 4.3 to higher dimensions $d\geq 2$ is to identify a transformation $F$ in $\mathbb{R}^{d}$ with analogous properties. However, preserving convex hulls is a much more restrictive condition in dimensions $d\geq 2$ that for instance forces $F$ to be affine if it is one-to-one: this is the so-called fundamental theorem of affine geometry, see e.g. Theorem 2.16 in Artstein-Avidan et al., (2012). One may hope that a weaker but still sufficient condition for the analysis would hold: namely, by inspecting the proof of Theorem 4.3, it can be noted that the condition
@@
would suffice. A natural candidate for $F$ is the Rosenblatt transformation, which maps continuous random vectors to vectors uniformly distributed on the hypercube by inductively projecting their coordinates with their conditional c.d.f. (Rosenblatt,, 1952). We show below that unfortunately, (8) does not hold for the Rosenblatt transformation.
@@ E.5 Center Separation Assumption when $\mathcal{E}=\mathcal{I}^{d}$ and $\mathcal{E}=\mathcal{S}^{d-1}$
When $\mathcal{E}=\mathbb{R}^{d}$, this assumption is reasonable: since the expected squared Euclidean distance between two uniformly sampled points scales linearly with $d$ (Anderssen et al.,, 1976), it is plausible for $\delta_{\min}^{2}$ to also scale with $d$. If $\mathcal{E}=\mathcal{I}^{d}$, it is reasonable if $\sigma$ is small. In this context, the $d$ in the numerator of the threshold formula
@@
Although Theorems B.13 and 4.5 still hold as stated for subgaussian mixtures supported on the sphere, a tailored analysis would employ concentration inequalities native to that geometry. They would require a weaker separability assumption and result in a different threshold rule, likely both without a linear dependence on $d$. Such an analysis is beyond the scope of this work. Our empirical evaluation explores this by treating the constant in the definition of the threshold $\hat{m}(t)$ as a tunable hyperparameter, so that we can find a practical balance suitable for each geometric setting.
@@ E.6 Challenges of the GHC Regret Analysis
where $M_{T}$ is the total number of incorrect guesses (mistakes). A mistake occurs if $q_{t}\in C_{k}$ for some $k\in[N]$ while the algorithm guesses $i\neq k$. It can be checked that this implies $d(q_{t},C_{i})\leq\varepsilon_{i,t}+\tau\varepsilon_{k,t}$, where $\varepsilon_{j,t}=d_{H}(\mathcal{C}_{j},\hat{\mathcal{C}}_{j,t})$ is the Hausdorff distance between the cell $\mathcal{C}_{j}$ and its estimate. Bounding these Hausdorff distances is already difficult due to the non-i.i.d. hull formation. Furthermore, in the simpler i.i.d. uniform case, bounds on the Hausdorff distance between a random polytope and the polytope it approximates scale as $n^{-1/d}$ where $n$ is the number of points (Bárány,, 1989) . This analysis approach would yield bounds on the number of mistakes scaling with $T^{(d-1)/d}$, which, asymptotically in $T$, is significantly worse than the $\log^{d}{T}$ bound of Theorem 4.1.
@@ F.1 Synthetic Experiments
F.1.1 Evaluation of CHC, GHC and Comparison with Sequential $k$-Means Baseline
First, we report the performance of GHC and CHC. We recall that CHC corresponds to GHC for $\tau=0.$ We also present the results of a baseline algorithm based on sequential $k$-means ($\operatorname{SKM}$) (MacQueen,, 1967). At each round, it stores $N$ estimated centroids $\{\hat{c}_{i}\}_{i\in[N]}$. If the algorithm guesses that the label of $q$ is $i$, the corresponding centroid is updated as $\hat{c}_{i}\leftarrow\hat{c}_{i}+(q-\hat{c}_{i})/\kappa_{i}$ where $\kappa_{i}$ is the number of times label $i$ was guessed up to the current round. This algorithm has two phases:
@@ F.1.1 Evaluation of CHC, GHC and Comparison with Sequential $k$-Means Baseline
    In phase 1, it calls the expert up until all labels are observed.
@@
    In phase 2, it guesses the label of the nearest centroid.
@@
Since this baseline never calls the expert after all labels are observed, it is a relevant comparison to GHC with threshold $\tau=1.$ As for GHC  we do not report the number of calls in the first phase, although they are included in the regret calculation, as we describe below.
@@
Results are reported in Tables 1 and 2 for $\mathcal{I}^{d}$ and $\mathcal{S}^{d-1}$ respectively. “C”, “M” and “R” denote the mean $\pm$ standard deviation of the number of expert calls in the second phase, the number of incorrect guesses and the cumulative regret respectively.
@@
Observe that the optimal threshold increases with the dimension: in dimension one, keeping a low threshold is optimal and the regret, mostly driven by the number of expert calls, is limited. Already in dimension $10$, CHC almost calls the expert for the $5,000$ queries, as expected from the discussion following Corollary 4.2. As the dimension grows large, the optimal threshold appears to rapidly approach one. In the mixture case, the performance of GHC is surprisingly better in large dimension. A likely explanation in the $\mathcal{E}=\mathcal{I}^{d}$ is the $\sqrt{d}$ scaling of the average distance between the uniformly sampled seeds in $\mathcal{I}^{d}$ (Anderssen et al.,, 1976), making the Gaussian clusters more separated since their covariance is fixed. The same regret improvement is observed when $\mathcal{E}=\mathcal{S}^{d-1}$, even though the average distance between the uniformly sampled seeds will remain bounded by $2$. This average distance in fact approaches $\sqrt{2}$ as $d\to\infty$, because the probability that two seeds are orthogonal approaches one as $d\to\infty$. This entails that the centers are more evenly separated in higher dimensions and could explain why better regret is observed in these cases. The performance of $\operatorname{SKM}$ is worse than GHC for any threshold in the uniform settings, and comparable to GHC with high threshold in the mixture settings.
Table 1: Performance metrics of GHC and Sequential K-Means ($\operatorname{SKM}$) on $\mathcal{I}^{d}$ ($T=5000$).
@@
Table 2: Performance metrics of GHC and Sequential K-Means ($\operatorname{SKM}$) on $\mathcal{S}^{d-1}$ ($T=5000$).
Dim Dist. $\tau=0.00$ $\tau=0.10$ $\tau=0.20$ $\tau=0.40$ $\tau=0.60$ $\tau=0.80$ $\tau=1.00$ SKM 2 Unif. C: $50\,\pm\,7$ M: 0± $0\,\pm\,$0 R: $132\,\pm\,12$ C: $43\,\pm\,5$ M: 1± $1\,\pm\,$1 R: $125\,\pm\,6$ C: $41\,\pm\,3$ M: $2\,\pm\,1$ R: $142\,\pm\,7$ C: $35\,\pm\,2$ M: $5\,\pm\,3$ R: $157\,\pm\,30$ C: $28\,\pm\,1$ M: $18\,\pm\,6$ R: $288\,\pm\,69$ C: 22± $2\,\pm\,2$ M: 44± $4\,\pm\,11$ R: $563\,\pm\,128$ C: 0± $0\,\pm\,$0 M: $315\,\pm\,148$ R: $3495\,\pm\,1633$ C: 0± $0\,\pm\,$0 M: $2431\,\pm\,174$ R: $26804\,\pm\,1923$ Mix. C: $61\,\pm\,3$ M: 0± $0\,\pm\,$0 R: $138\,\pm\,8$ C: $47\,\pm\,4$ M: 0± $0\,\pm\,$0 R: $114\,\pm\,8$ C: $38\,\pm\,2$ M: 1± $1\,\pm\,$1 R: $104\,\pm\,13$ C: $29\,\pm\,1$ M: $3\,\pm\,1$ R: $110\,\pm\,17$ C: $24\,\pm\,1$ M: $9\,\pm\,3$ R: $166\,\pm\,31$ C: $18\,\pm\,2$ M: 22± $2\,\pm\,9$ R: $299\,\pm\,98$ C: 0± $0\,\pm\,$0 M: $345\,\pm\,77$ R: $3810\,\pm\,847$ C: 0± $0\,\pm\,$0 M: $70\,\pm\,58$ R: $791\,\pm\,645$ 4 Unif. C: $602\,\pm\,15$ M: 0± $0\,\pm\,$0 R: $1225\,\pm\,31$ C: $514\,\pm\,11$ M: 0± $0\,\pm\,$0 R: $1053\,\pm\,26$ C: $438\,\pm\,11$ M: $2\,\pm\,1$ R: $922\,\pm\,25$ C: $329\,\pm\,10$ M: $14\,\pm\,3$ R: $836\,\pm\,40$ C: $246\,\pm\,6$ M: 55± $5\,\pm\,11$ R: $1122\,\pm\,108$ C: $177\,\pm\,7$ M: $137\,\pm\,4$ R: $1878\,\pm\,52$ C: 0± $0\,\pm\,$0 M: $1909\,\pm\,339$ R: $21016\,\pm\,3728$ C: 0± $0\,\pm\,$0 M: $2329\,\pm\,742$ R: $25645\,\pm\,8157$ Mix. C: $531\,\pm\,29$ M: 0± $0\,\pm\,$0 R: $1083\,\pm\,56$ C: $127\,\pm\,4$ M: 0± $0\,\pm\,$0 R: $275\,\pm\,5$ C: $51\,\pm\,4$ M: 0± $0\,\pm\,$0 R: $124\,\pm\,11$ C: $12\,\pm\,2$ M: 0± $0\,\pm\,$0 R: $46\,\pm\,7$ C: $3\,\pm\,2$ M: 0± $0\,\pm\,$0 R: $28\,\pm\,3$ C: 0± $0\,\pm\,$0 M: 0± $0\,\pm\,$0 R: 22± $2\,\pm\,6$ C: 0± $0\,\pm\,$0 M: 0± $0\,\pm\,$0 R: 22± $2\,\pm\,7$ C: 0± $0\,\pm\,$0 M: 0± $0\,\pm\,$0 R: 22± $2\,\pm\,7$ 10 Unif. C: $2428\,\pm\,72$ M: 0± $0\,\pm\,$0 R: $4878\,\pm\,150$ C: $2167\,\pm\,74$ M: 0± $0\,\pm\,$0 R: $4357\,\pm\,154$ C: 1919± $19\,\pm\,74$ M: 0± $0\,\pm\,$0 R: $3862\,\pm\,155$ C: $1430\,\pm\,54$ M: $6\,\pm\,3$ R: $2952\,\pm\,138$ C: $1011\,\pm\,42$ M: $46\,\pm\,6$ R: $2550\,\pm\,123$ C: $632\,\pm\,21$ M: $194\,\pm\,21$ R: $3423\,\pm\,225$ C: 0± $0\,\pm\,$0 M: $2990\,\pm\,94$ R: $32914\,\pm\,1023$ C: 0± $0\,\pm\,$0 M: $3630\,\pm\,358$ R: $39956\,\pm\,3931$ Mix. C: $3735\,\pm\,38$ M: 0± $0\,\pm\,$0 R: $7490\,\pm\,74$ C: $1439\,\pm\,36$ M: 0± $0\,\pm\,$0 R: $2898\,\pm\,63$ C: $430\,\pm\,13$ M: 0± $0\,\pm\,$0 R: $879\,\pm\,21$ C: 55± $5\,\pm\,3$ M: 0± $0\,\pm\,$0 R: $130\,\pm\,7$ C: $6\,\pm\,3$ M: 0± $0\,\pm\,$0 R: $31\,\pm\,7$ C: 0± $0\,\pm\,$0 M: 0± $0\,\pm\,$0 R: $20\,\pm\,11$ C: 0± $0\,\pm\,$0 M: 0± $0\,\pm\,$0 R: $20\,\pm\,11$ C: 0± $0\,\pm\,$0 M: 0± $0\,\pm\,$0 R: $20\,\pm\,11$ 50 Unif. C: $4979\,\pm\,5$ M: 0± $0\,\pm\,$0 R: $9983\,\pm\,4$ C: $4964\,\pm\,6$ M: 0± $0\,\pm\,$0 R: $9953\,\pm\,6$ C: $4936\,\pm\,7$ M: 0± $0\,\pm\,$0 R: 9898± $98\,\pm\,8$ C: $4720\,\pm\,9$ M: 0± $0\,\pm\,$0 R: $9464\,\pm\,18$ C: $4119\,\pm\,20$ M: 0± $0\,\pm\,$0 R: $8262\,\pm\,41$ C: $2825\,\pm\,13$ M: $42\,\pm\,8$ R: $6135\,\pm\,73$ C: 0± $0\,\pm\,$0 M: $3502\,\pm\,53$ R: $38549\,\pm\,582$ C: 0± $0\,\pm\,$0 M: $3747\,\pm\,99$ R: $41246\,\pm\,1091$ Mix. C: $4990\,\pm\,2$ M: 0± $0\,\pm\,$0 R: $10000\,\pm\,0$ C: $4990\,\pm\,2$ M: 0± $0\,\pm\,$0 R: $10000\,\pm\,0$ C: $4971\,\pm\,4$ M: 0± $0\,\pm\,$0 R: $9961\,\pm\,12$ C: $1650\,\pm\,28$ M: 0± $0\,\pm\,$0 R: $3319\,\pm\,56$ C: $109\,\pm\,7$ M: 0± $0\,\pm\,$0 R: $237\,\pm\,15$ C: $5\,\pm\,2$ M: 0± $0\,\pm\,$0 R: $28\,\pm\,5$ C: 0± $0\,\pm\,$0 M: 0± $0\,\pm\,$0 R: $19\,\pm\,5$ C: 0± $0\,\pm\,$0 M: 0± $0\,\pm\,$0 R: $19\,\pm\,5$

### v2

@@ E.2 Asymptotic Lower bound on the Regret of CHC
Now for any $0<a\leq 1$, $\lambda(\{y\in[0,1]^{d},\prod y_{i}\leq a\})=a\sum_{k=0}^{d-1}\log(1/a)^{k}/k!$ (see Bárány and Larman (1988), p.288). Consequently, $w(t)\leq 2t\sum_{k=0}^{d-1}\log(1/(2t))^{k}/k!$ whenever $0<t\leq 1/2$. Applying this to $t=1/n$ yields the result. ∎
When $\nu$ is uniform on some arbitrary polytope $P$, explicitly bounding the volume of the wet part of $P$ becomes significantly more difficult. Schütt (1991) proved in their Lemma 1.8 that for some $t_{d}$ and $c_{d}$ that only depends on $d$, if $0<t<t_{d}$,
@@
A natural idea to generalize Theorem 3.3 to higher dimensions $d\geq 2$ is to identify a transformation $F$ in $\mathbb{R}^{d}$ with analogous properties. However, preserving convex hulls is a much more restrictive condition in dimensions $d\geq 2$ that for instance forces $F$ to be affine if it is one-to-one: this is the so-called fundamental theorem of affine geometry, see e.g. Theorem 2.16 in Artstein-Avidan et al. (2012). One may hope that a weaker but still sufficient condition for the analysis would hold: namely, by inspecting the proof of Theorem 3.3, it can be noted that the condition
@@
would suffice. A natural candidate for $F$ is the Rosenblatt transformation, which maps continuous random vectors to vectors uniformly distributed on the hypercube by inductively projecting their coordinates with their conditional c.d.f. (Rosenblatt, 1952). We show below that unfortunately, (8) does not hold for the Rosenblatt transformation.
@@ E.5 Center Separation Assumption when $\mathcal{E}=\mathcal{I}^{d}$ and $\mathcal{E}=\mathcal{S}^{d-1}$
When $\mathcal{E}=\mathbb{R}^{d}$, this assumption is reasonable: since the expected squared Euclidean distance between two uniformly sampled points scales linearly with $d$ (Anderssen et al., 1976), it is plausible for $\delta_{\min}^{2}$ to also scale with $d$. If $\mathcal{E}=\mathcal{I}^{d}$, it is reasonable if $\sigma$ is small. In this context, the $d$ in the numerator of the threshold formula
@@
Although Theorems B.13 and 3.5 still hold as stated for subgaussian mixtures supported on the sphere, a tailored analysis would employ concentration inequalities native to that geometry. They would require a weaker separability assumption and result in a different threshold rule, likely both without a linear dependence on $d$. Such an analysis is beyond the scope of this work. Our empirical evaluation explores this by treating the constant in the definition of the threshold $\hat{m}(t)$ as a tunable hyperparameter, so that we can find a practical balance suitable for each geometric setting.
@@ E.6 Challenges of the GHC Regret Analysis
where $M_{T}$ is the total number of incorrect guesses (mistakes). A mistake occurs if $q_{t}\in C_{k}$ for some $k\in[N]$ while the algorithm guesses $i\neq k$. It can be checked that this implies $d(q_{t},C_{i})\leq\varepsilon_{i,t}+\tau\varepsilon_{k,t}$, where $\varepsilon_{j,t}=d_{H}(\mathcal{C}_{j},\hat{\mathcal{C}}_{j,t})$ is the Hausdorff distance between the cell $\mathcal{C}_{j}$ and its estimate. Bounding these Hausdorff distances is already difficult due to the non-i.i.d. hull formation. Furthermore, in the simpler i.i.d. uniform case, bounds on the Hausdorff distance between a random polytope and the polytope it approximates scale as $n^{-1/d}$ where $n$ is the number of points (Bárány, 1989) . This analysis approach would yield bounds on the number of mistakes scaling with $T^{(d-1)/d}$, which, asymptotically in $T$, is significantly worse than the $\log^{d}{T}$ bound of Theorem 3.1.
@@ F.1 Synthetic Experiments
F.1.1 Evaluation of CHC, GHC and Comparison with Baselines
First, we report the performance of GHC and CHC. We recall that CHC corresponds to GHC for $\tau=0.$ We also present the results of two baseline algorithms.
Sequential $k$-Means (SKM) baseline.
The first baseline is based on sequential $k$-means (MacQueen, 1967). At each round, it stores $N$ estimated centroids $\{\hat{c}_{i}\}_{i\in[N]}$. If the algorithm guesses that the label of $q$ is $i$, the corresponding centroid is updated as $\hat{c}_{i}\leftarrow\hat{c}_{i}+(q-\hat{c}_{i})/\kappa_{i}$ where $\kappa_{i}$ is the number of times label $i$ was guessed up to the current round. This algorithm has two phases:
@@ F.1.1 Evaluation of CHC, GHC and Comparison with Baselines
    In phase 1 (initialization phase), the expert is called until all labels are observed.
@@
    In phase 2 (unsupervised phase), the algorithm predicts the label of the nearest centroid: $\hat{\imath}=\arg\min_{i\in[N]}\|q-\hat{c}_{i}\|_{2}$
@@
Since this baseline never calls the expert after all labels are observed, it is a relevant comparison to GHC with threshold $\tau=1.$
Active Multiclass Perceptron (AMP) baseline.
The second baseline is a selective variant of the multiclass perceptron. The standard multiclass perceptron was studied in Crammer and Singer (2003), and in the binary classification setting, an active variant was proposed in Cesa-bianchi et al. (2004). Although these works study an adversarial setting under a margin assumption, whereas we study the stochastic i.i.d. regime, such a method is a reasonable baseline in practice if the classes are separated by hyperplanes passing through the origin. This is the case when the query domain is $\mathcal{S}^{d-1}$.
The AMP baseline takes as input a threshold $\tau\in[0,1]$ and first initializes weights $w_{1},\dots,w_{N}$ at 0. Then at each round $t=1,\dots,T$:
  1. (i)
    It calculates normalized scores $q_{t}^{\top}v_{i}$ where $v_{i}=w_{i}/\|w_{i}\|$ for the query $q_{t}$ and identifies the best candidate label $j=\arg\max_{i}q_{t}^{\top}v_{i}$ and second-best candidate label $k=\arg\max_{i\neq j}q_{t}^{\top}v_{j}$.
  2. (ii)
    If $q_{t}^{\top}(v_{j}-v_{k})\leq 2(1-\tau)$, it queries the expert for the true label $i_{t}$. It then updates weights by setting $w_{j}\leftarrow w_{j}+q_{t}$ if $j=i_{t}$, and $w_{j}\leftarrow w_{j}-q_{t}$ if $j\neq i_{t}$.
  3. (iii)
    If $q_{t}^{\top}(v_{j}-v_{k})>2(1-\tau)$, it predicts $\hat{\imath}_{t}=j$.
We note that typically, Perceptron (and its active variants) only update weights when the prediction is wrong (Cesa-bianchi et al., 2004), i.e. it would set $w_{i_{t}}\leftarrow w_{i_{t}}+q_{t}$ and $w_{j}\leftarrow w_{j}-q_{t}$ only if $j\neq i_{t}$. However, we found better empirical performance by also updating weights when the prediction matches the true label.
@@
For each setting, we report the performance metrics of AMP on the sphere for the threshold value with lowest regret among $\{0.05k:k\in\{0,\dots,20\}\}$. As for GHC, we do not report the number of calls in the first phase, although they are included in the regret calculation. Results are reported in Tables 1 and 2 for $\mathcal{I}^{d}$ and $\mathcal{S}^{d-1}$ respectively. “C”, “M” and “R” denote the mean $\pm$ standard deviation of the number of expert calls in the second phase, the number of incorrect guesses and the cumulative regret respectively.
@@
We observe that the optimal threshold increases with the dimension: in dimension one, keeping a low threshold is optimal and the regret, mostly driven by the number of expert calls, is limited. Already in dimension $10$, CHC almost calls the expert for the $5,000$ queries, as expected from the discussion following Corollary 3.2. As the dimension grows large, the optimal threshold appears to rapidly approach one. In the mixture case, the performance of GHC is surprisingly better in large dimension. A likely explanation in the $\mathcal{E}=\mathcal{I}^{d}$ is the $\sqrt{d}$ scaling of the average distance between the uniformly sampled seeds in $\mathcal{I}^{d}$ (Anderssen et al., 1976), making the Gaussian clusters more separated since their covariance is fixed. The same regret improvement is observed when $\mathcal{E}=\mathcal{S}^{d-1}$, even though the average distance between the uniformly sampled seeds will remain bounded by $2$. This average distance in fact approaches $\sqrt{2}$ as $d\to\infty$, because the probability that two seeds are orthogonal approaches one as $d\to\infty$. This entails that the centers are more evenly separated in higher dimensions and could explain why better regret is observed in these cases. The performance of SKM is worse than GHC for any threshold in the uniform settings, and comparable to GHC with high threshold in the mixture settings. GHC also always outperforms the AMP baseline, especially in lower dimensions.
Table 1: Performance metrics of GHC and SKM on $\mathcal{I}^{d}$ ($T=5000$).
@@
Table 2: Performance metrics of GHC, SKM and AMP on $\mathcal{S}^{d-1}$ ($T=5000$).
Dim Dist. $\tau=0.00$ $\tau=0.10$ $\tau=0.20$ $\tau=0.40$ $\tau=0.60$ $\tau=0.80$ $\tau=1.00$ SKM AMP 2 Unif. C: $50\,\pm\,7$ M: 0± $0\,\pm\,$0 R: $132\,\pm\,12$ C: $43\,\pm\,5$ M: 1± $1\,\pm\,$1 R: $125\,\pm\,6$ C: $41\,\pm\,3$ M: $2\,\pm\,1$ R: $142\,\pm\,7$ C: $35\,\pm\,2$ M: $5\,\pm\,3$ R: $157\,\pm\,30$ C: $28\,\pm\,1$ M: $18\,\pm\,6$ R: $288\,\pm\,69$ C: 22± $2\,\pm\,2$ M: 44± $4\,\pm\,11$ R: $563\,\pm\,128$ C: 0± $0\,\pm\,$0 M: $315\,\pm\,148$ R: $3495\,\pm\,1633$ C: 0± $0\,\pm\,$0 M: $2431\,\pm\,174$ R: $26804\,\pm\,1923$ C: $1856\,\pm\,46$ M: $114\,\pm\,42$ R: $4967\,\pm\,421$ Mix. C: $61\,\pm\,3$ M: 0± $0\,\pm\,$0 R: $138\,\pm\,8$ C: $47\,\pm\,4$ M: 0± $0\,\pm\,$0 R: $114\,\pm\,8$ C: $38\,\pm\,2$ M: 1± $1\,\pm\,$1 R: $104\,\pm\,13$ C: $29\,\pm\,1$ M: $3\,\pm\,1$ R: $110\,\pm\,17$ C: $24\,\pm\,1$ M: $9\,\pm\,3$ R: $166\,\pm\,31$ C: $18\,\pm\,2$ M: 22± $2\,\pm\,9$ R: $299\,\pm\,98$ C: 0± $0\,\pm\,$0 M: $345\,\pm\,77$ R: $3810\,\pm\,847$ C: 0± $0\,\pm\,$0 M: $70\,\pm\,58$ R: $791\,\pm\,645$ C: $5000\,\pm\,0$ M: 0± $0\,\pm\,$0 R: $10000\,\pm\,0$ 4 Unif. C: $602\,\pm\,15$ M: 0± $0\,\pm\,$0 R: $1225\,\pm\,31$ C: $514\,\pm\,11$ M: 0± $0\,\pm\,$0 R: $1053\,\pm\,26$ C: $438\,\pm\,11$ M: $2\,\pm\,1$ R: $922\,\pm\,25$ C: $329\,\pm\,10$ M: $14\,\pm\,3$ R: $836\,\pm\,40$ C: $246\,\pm\,6$ M: 55± $5\,\pm\,11$ R: $1122\,\pm\,108$ C: $177\,\pm\,7$ M: $137\,\pm\,4$ R: $1878\,\pm\,52$ C: 0± $0\,\pm\,$0 M: $1909\,\pm\,339$ R: $21016\,\pm\,3728$ C: 0± $0\,\pm\,$0 M: $2329\,\pm\,742$ R: $25645\,\pm\,8157$ C: $869\,\pm\,23$ M: $96\,\pm\,10$ R: $2799\,\pm\,110$ Mix. C: $531\,\pm\,29$ M: 0± $0\,\pm\,$0 R: $1083\,\pm\,56$ C: $127\,\pm\,4$ M: 0± $0\,\pm\,$0 R: $275\,\pm\,5$ C: $51\,\pm\,4$ M: 0± $0\,\pm\,$0 R: $124\,\pm\,11$ C: $12\,\pm\,2$ M: 0± $0\,\pm\,$0 R: $46\,\pm\,7$ C: $3\,\pm\,2$ M: 0± $0\,\pm\,$0 R: $28\,\pm\,3$ C: 0± $0\,\pm\,$0 M: 0± $0\,\pm\,$0 R: 22± $2\,\pm\,6$ C: 0± $0\,\pm\,$0 M: 0± $0\,\pm\,$0 R: 22± $2\,\pm\,7$ C: 0± $0\,\pm\,$0 M: 0± $0\,\pm\,$0 R: 22± $2\,\pm\,7$ C: $70\,\pm\,46$ M: $16\,\pm\,8$ R: $312\,\pm\,170$ 10 Unif. C: $2428\,\pm\,72$ M: 0± $0\,\pm\,$0 R: $4878\,\pm\,150$ C: $2167\,\pm\,74$ M: 0± $0\,\pm\,$0 R: $4357\,\pm\,154$ C: 1919± $19\,\pm\,74$ M: 0± $0\,\pm\,$0 R: $3862\,\pm\,155$ C: $1430\,\pm\,54$ M: $6\,\pm\,3$ R: $2952\,\pm\,138$ C: $1011\,\pm\,42$ M: $46\,\pm\,6$ R: $2550\,\pm\,123$ C: $632\,\pm\,21$ M: $194\,\pm\,21$ R: $3423\,\pm\,225$ C: 0± $0\,\pm\,$0 M: $2990\,\pm\,94$ R: $32914\,\pm\,1023$ C: 0± $0\,\pm\,$0 M: $3630\,\pm\,358$ R: $39956\,\pm\,3931$ C: $1415\,\pm\,30$ M: $97\,\pm\,14$ R: $3902\,\pm\,158$ Mix. C: $3735\,\pm\,38$ M: 0± $0\,\pm\,$0 R: $7490\,\pm\,74$ C: $1439\,\pm\,36$ M: 0± $0\,\pm\,$0 R: $2898\,\pm\,63$ C: $430\,\pm\,13$ M: 0± $0\,\pm\,$0 R: $879\,\pm\,21$ C: 55± $5\,\pm\,3$ M: 0± $0\,\pm\,$0 R: $130\,\pm\,7$ C: $6\,\pm\,3$ M: 0± $0\,\pm\,$0 R: $31\,\pm\,7$ C: 0± $0\,\pm\,$0 M: 0± $0\,\pm\,$0 R: $20\,\pm\,11$ C: 0± $0\,\pm\,$0 M: 0± $0\,\pm\,$0 R: $20\,\pm\,11$ C: 0± $0\,\pm\,$0 M: 0± $0\,\pm\,$0 R: $20\,\pm\,11$ C: $23\,\pm\,9$ M: 0± $0\,\pm\,$0 R: 50± $\pm\,21 50$ Unif. C: $4979\,\pm\,5$ M: 0± $0\,\pm\,$0 R: $9983\,\pm\,4$ C: $4964\,\pm\,6$ M: 0± $0\,\pm\,$0 R: $9953\,\pm\,6$ C: $4936\,\pm\,7$ M: 0± $0\,\pm\,$0 R: 9898± $98\,\pm\,8$ C: $4720\,\pm\,9$ M: 0± $0\,\pm\,$0 R: $9464\,\pm\,18$ C: $4119\,\pm\,20$ M: 0± $0\,\pm\,$0 R: $8262\,\pm\,41$ C: $2825\,\pm\,13$ M: $42\,\pm\,8$ R: $6135\,\pm\,73$ C: 0± $0\,\pm\,$0 M: $3502\,\pm\,53$ R: $38549\,\pm\,582$ C: 0± $0\,\pm\,$0 M: $3747\,\pm\,99$ R: $41246\,\pm\,1091$ C: $2879\,\pm\,32$ M: $94\,\pm\,17$ R: $6796\,\pm\,156$ Mix. C: $4990\,\pm\,2$ M: 0± $0\,\pm\,$0 R: $10000\,\pm\,0$ C: $4990\,\pm\,2$ M: 0± $0\,\pm\,$0 R: $10000\,\pm\,0$ C: $4971\,\pm\,4$ M: 0± $0\,\pm\,$0 R: $9961\,\pm\,12$ C: $1650\,\pm\,28$ M: 0± $0\,\pm\,$0 R: $3319\,\pm\,56$ C: $109\,\pm\,7$ M: 0± $0\,\pm\,$0 R: $237\,\pm\,15$ C: $5\,\pm\,2$ M: 0± $0\,\pm\,$0 R: $28\,\pm\,5$ C: 0± $0\,\pm\,$0 M: 0± $0\,\pm\,$0 R: $19\,\pm\,5$ C: 0± $0\,\pm\,$0 M: 0± $0\,\pm\,$0 R: $19\,\pm\,5$ C: $10\,\pm\,0$ M: 1± $1\,\pm\,$1 R: $28\,\pm\,8$